how do i solve 12=3/(x+4)
Found 2 solutions by nerdybill, MathLover1:
Answer by nerdybill(7384)

12=3/(x+4)
multiply both sides by (x+4) to get:
12(x+4) = 3
distribute the 12:
12x+48 = 3
subtract 48 from both sides:
12x = -45
x = -45/12
reduce fraction:
x = -15/4

Answer by MathLover1(20850)